OpenStudy (anonymous):

A bag contains 26 Scrabble tiles, one for each letter of the alphabet. We draw 2 tiles from the bag at random, we do replace each tile after it is drawn. a) How many ways are there to draw two vowels, for example AA, AE, EA, b) What is the probability of drawing two vowels? c) What is the probability of drawing two consonants, for example BC, CB, BD d) What is the probability of drawing one vowel and one consonant, for example AB, BA, AC? e) What is the sum of your answers to parts b), c), d), and e) of this question?

OpenStudy (shiraz14):

Number of vowels = 5 a) Number of ways to draw 2 vowels = (5 x 5)-5 = 20 b) P(drawing 2 vowels with replacement) = (5/26)^2 = 25/676 c) P(drawing 2 consonants) = (21/26)^2 = 441/676 d) P(drawing 1 vowel & 1 consonant) = (5/26 x 21/26) X 2 = 210/676 e) Sum = 1

OpenStudy (anonymous):

thankyou very much!!

OpenStudy (shiraz14):

You're welcome & a medal would be much appreciated.